Find 3 consecutive odd integers such that 4 less than twice the second has the same value as 3 times the number that is one more than the third
Answer by stanbon(75887)

Find 3 consecutive odd integers such that 4 less than twice the second has the same value as 3 times the number that is one more than the third
----
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
------
Equation:
2(2x+1)-4 = 3(2x+4)
----
4x-2 = 6x+12
2x = -14
x = -7
---------------
1st: 2x-1 = -15
2nd: 2x+1 = -13
3rd: 2x+3 = -11
======================
Cheers,
Stan H.